# Composing Codensity Bisimulations

### v1

@@ 3.2 Composing $F$-Coalgebras
Let $F:\mathbb{C}\to\mathbb{C}$ be an endofunctor over a category $\mathbb{C}$. We model an $N$-ary composition operation for $F$-coalgebras by a pair of
@@
    a functor $T:\mathbb{C}^{N}\to\mathbb{C}$ (called structure functor) describing how the operation transforms coalgebra carriers, and
@@
The second component induces the lifting ${T}_{\lambda}:(\mathbf{Coalg}(F))^{N}\to\mathbf{Coalg}(F)$ of $T$ along $(U_{F})^{N},U_{F}$ given by
${T}_{\lambda}\triangleq\lambda\circ T(c_{1},\cdots,c_{N}).$
@@
We introduce a binary one-step composition operation (×$(\times,\lambda)$ on ${\mathcal{P}}$-coalgebras. It takes the binary product of coalgebra carriers. The distributive law $\lambda^{{\mathcal{P}}}_{X,Y}:{\mathcal{P}}X\times{\mathcal{P}}Y\to{\mathcal{P}}(X\times Y)$ is given by $\lambda^{{\mathcal{P}}}_{X,Y}(A,B)\triangleq A\times B$.
@@
$\lambda^{\mathrm{da}}_{X,Y}((t_{1},\rho_{1}),(t_{2},\rho_{2}))\triangleq(t_{1}\land t_{2},~{}\lambda a~{}.~{}(\rho_{1}(a),\rho_{2}(a))).$
@@ 3.3 Composing Coalgebraic Bismulations
Let $(F,{\dot{F}}):p\to p$ and $(T,{\dot{T}}):p^{N}\to p$ be 1-cells in $\mathbf{CAT}^{\mathbf{CLat}_{\sqcap}}$ and $(\lambda,\dot{\lambda}):(T,{\dot{T}})\circ(F^{N},{\dot{F}}^{N})\to(F,{\dot{F}})\circ(T,{\dot{T}})$ be a 2-cell. Then $(T_{\lambda},\dot{T}_{\dot{\lambda}}):(\mathbf{Coalg}(p))^{N}\to\mathbf{Coalg}(p)$ is a 1-cell in $\mathbf{CAT}^{\to}$.∎
A restricted version of this theorem is also available. Let $C$ be a subclass of objects in $\mathbb{E}$. If $\lambda$ is liftable on $C^{N}$, then we obtain a 1-cell $(T_{\lambda},\dot{T}_{\dot{\lambda}}):(\mathbf{Coalg}(p)|_{C})^{N}\to\mathbf{Coalg}(p)$; here $\mathbf{Coalg}(p)|_{C}$ is the restriction of $\mathbf{Coalg}(p)$ to the category of ${\dot{F}}$-coalgebras over objects in $C$. This restriction will be used after Proposition 7.3.
@@
In the setting of Theorem 3.9, for any $F$-coalgebra $c_{i}:X_{i}\to F(X_{i})$ ($i=1\cdots N$),
@@
Let $(T,\lambda)$ be an $N$-ary one-step composition operation for $F$-coalgebras. 1. How do we non-trivially lift structure functor (such as $N$-ary product functors) in order to capture composition operators at the level of relations? 2. When is $\lambda$ liftable as a distributive law between codensity liftings of $F$ and $T$?
To answer to these questions, we employ Beohar et al.’s recent decomposition result of codensity liftings [3]. Their decomposition separates fibration-specific parts in the codensity lifting from the central part that does the actual lifting task. Upon their decomposition, in Section 5, we first answer to (1) by generalizing the codensity lifting to arbitrary functors. Then in Section 6, we answer to (2) by providing a sufficient condition to lift $\lambda$ to a distributive law between codensity liftings. The proof of the sufficiency takes advantage of the decomposition and a 2-categorical nature of the central part in the decomposition.
@@ 4 Generalizing the Codensity Lifting
We recall the multiple parameter codensity lifting in [29]. It lifts an endofunctor $F:\mathbb{B}\to\mathbb{B}$ along a ${\mathbf{CLat}_{\sqcap}}$-fibration $p:\mathbb{E}\to\mathbb{B}$ using three parameters $\Omega,\mathbf{\Omega},\tau$. We pack $(p,\Omega,\mathbf{\Omega})$ into the following data.
@@
Let $A$ be a discrete category. A ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values is a tuple $(p:\mathbb{E}\to\mathbb{B},\Omega:A\to\mathbb{B},\mathbf{\Omega}:A\to\mathbb{E})$ of functors such that $\Omega=p\circ\mathbf{\Omega}$ and $p$ is a ${\mathbf{CLat}_{\sqcap}}$-fibration. The notation for a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values is $(\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbb{E}\to\mathbb{B})$.
Let $(\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbb{E}\to\mathbb{B})$ be a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values. The codensity lifting of a functor $F:\mathbb{B}\to\mathbb{B}$ takes as argument a natural transformation $\tau:F\circ\Omega\to\Omega$. The components $\tau_{a}\colon F\circ\Omega(a)\to\Omega(a)$ are referred to in coalgebraic modal logic as modalities; $\tau$ is just an indexed collection of these, and below we will often refer to $\tau$ itself as a modality. The codensity lifting of $F:\mathbb{B}\to\mathbb{B}$ with $\tau$ along $p$ is given by
$[F,\tau]^{\mathbf{\Omega}}X\triangleq\bigsqcap_{a\in A,k\in\mathbb{E}(X,\mathbf{\Omega}(a))}(\tau_{a}\circ F(pk))^{*}(\mathbf{\Omega}(a)).$
It is an endofunctor over $\mathbb{E}$, and is a lifting of $F$ along $p$. When $A=1$, it is called the single-parameter codensity lifting. ${\dot{F}}$-bisimulations whose lifting part ${\dot{F}}$ is a codensity lifting are called codensity bisimulation.
@@
We write $2$ for the two-point set $\{\mathrm{false},\mathrm{true}\}$ and $\mathrm{Eq}_{2}$ for the object $(2,\{(x,x)~{}|~{}x\in 2\})$ in $\mathbf{EqRel}$. We identify it as a functor of type $1\to\mathbf{EqRel}$, and form a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $(2,\mathrm{Eq}_{2}):Id_{1}\to U_{\mathbf{EqRel}}$. The codensity lifting of ${\mathcal{P}}$ with the modality $\diamond:{\mathcal{P}}(2)\to 2$ given by $\diamond(U)=true\iff true\in U$ coincides with the lifting $\dot{\mathcal{P}}$ in Example 3.3 [23]. Therefore the standard bisimulations on ${\mathcal{P}}$-coalgebras are expressible as codensity bisimulations.
@@
We adopt the following ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values   $(\Omega,\mathbf{\Omega}):Id_{\Sigma\uplus\{\epsilon\}}\to U_{\mathbf{EqRel}}$ where $\Omega$ and $\mathbf{\Omega}$ are functors constantly returning $2$ and $\mathrm{Eq}_{2}$, respectively. We define the modality for ${F_{\mathrm{da}}}$ by
@@
Codensity bisimulation data consists of a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $(\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbb{E}\to\mathbb{B})$, an endofunctor $F:\mathbb{B}\to\mathbb{B}$, and a natural transformation $\tau:F\circ\Omega\to\Omega$.
@@ 4.1 Decomposition of Codensity Lifting
Let $((\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbb{E}\to\mathbb{B}),F,\tau)$ be codensity bisimulation data. Recently, Beohar et al. [3] introduced a decomposition of the codensity lifting as a “sandwich” of a monotone function within Galois connections. We adopt their decomposition as it is useful for analyzing the interaction between codensity liftings. In this paper, we describe their decomposition in fibered category theory. The decomposition is given as
@@
The description of the components of (2) is in order. First, we observe that the product $[A,s^{op}]\colon[A,\mathbf{Pred}^{op}]\to[A,\mathbf{Set}^{op}]$ of $A$-fold copies of the opposite of the subobject fibration $U_{\mathbf{Pred}}:\mathbf{Pred}\to\mathbf{Set}$ is a ${\mathbf{CLat}_{\sqcap}}$-fibration. Now the adjunction $L^{p,\mathbf{\Omega}}\dashv R^{p,\mathbf{\Omega}}$ in (2) is given between $\mathbb{E}$ and the category obtained by the change-of-base of the fibration $[A,U_{\mathbf{Pred}}^{op}]$ along the functor $H(b)\triangleq(\mathbb{B}(b,\Omega(-)))^{op}$:
@@
Concretely, an object of $\mathrm{Sp}^{A}(\mathbb{B},\Omega)$ is a pair of an object $X\in\mathbb{B}$ and a mapping $S:A\to\mathbf{Set}$ such that for any $a\in A$, $S(a)\subseteq\mathbb{B}(X,\Omega(a))$. A morphism from $(X,S)$ to $(Y,T)$ is a morphism $f:X\to Y$ such that for any $a\in A$ and $k\in T(a)$, $k\circ f\in S(a)$. The evident projection functor $r^{A}_{\mathbb{B},\Omega}:\mathrm{Sp}^{A}(\mathbb{B},\Omega)\to\mathbb{B}$ is a ${\mathbf{CLat}_{\sqcap}}$-fibration, because $s$ is so.
@@
$L^{p,\mathbf{\Omega}}(P)$ $\triangleq(pP,\lambda a~{}.~{}p(\mathbb{E}(P,\mathbf{\Omega}(a)))),$ $L^{p,\mathbf{\Omega}}(f)\triangleq f,$
@@
The middle part $\mathrm{Sp}^{A}(\mathbb{B},\Omega)$ in (2) is defined as a lifting of an endofunctor $F:\mathbb{B}\to\mathbb{B}$ along $r^{A}_{\mathbb{B},\Omega}$ using a modality $\tau:F\circ\Omega\to\Omega$. It is an extension of the predicate lifting $(P:X\to\Omega)\mapsto(\tau\circ F(P):F(X)\to\Omega)$, which is commonly used in the theory of coalgebras, to sets of morphisms into $\Omega$.
$\mathrm{Sp}^{A}(F,\tau)(X,S)$ $\triangleq(FX,\lambda a~{}.~{}\{\tau_{a}\circ Fk~{}|~{}k\in S(a)\})$ (3)
$\mathrm{Sp}^{A}(F,\tau)(f)$ $\triangleq Ff.$ (4)
@@ 4.2 Generalizing Codensity Liftings
  • A 0-cell is a pair of a category $\mathbb{C}$ with a functor $\Omega:A\to\mathbb{C}$.
  • A 1-cell from $(\mathbb{C},\Omega)$ to $(\mathbb{D},\Pi)$ is a pair of a functor $F:\mathbb{C}\to\mathbb{D}$ and a natural transformation $\tau:F\circ\Omega\to\Pi$.
  • A 2-cell from $(F,\tau):(\mathbb{C},\Omega)\to(\mathbb{D},\Pi)$ to $(G,\upsilon)$ is a natural transformation $\alpha:F\to G$ such that $\upsilon\mathbin{\bullet}(\alpha\circ\Omega)=\tau$.
Observe that an endo-1-cell on $(\mathbb{C},\Omega)$ bijectively corresponds to a pair of an endofunctor $F$ on $\mathbb{C}$ and a modality $\tau:F\circ\Omega\to\Omega$.
@@
Here, the functor $\mathrm{Sp}^{A}(F,\tau)$ is the extension of (3) to arbitrary 1-cell $(F,\tau):(\mathbb{C},\Omega)\to(\mathbb{D},\Pi)$ by the same formulas (3),(4). Moreover, $\mathrm{Sp}^{A}(F,\tau)$ preserves Cartesian morphisms and all meets. ∎
@@
Let $(\Omega,\mathbf{\Omega}):Id_{A}\to(p:\mathbb{E}\to\mathbb{B})$ and $(\Pi,\mathbf{\Pi}):Id_{A}\to(q:\mathbb{F}\to\mathbb{C})$ be ${\mathbf{CLat}_{\sqcap}}$-fibrations with truth values. The codensity lifting of $F:\mathbb{B}\to\mathbb{C}$ along $p,q$ with a natural transformation $\tau:F\circ\Omega\to\Pi$, denoted by $[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}$, is defined by
$[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}$ $\triangleq R^{q,\mathbf{\Pi}}\circ\mathrm{Sp}^{A}(F,\tau)\circ L^{p,\mathbf{\Omega}}$ (5)
@@
For any $(F,\dot{F})\in\mathbf{CAT}^{\mathbf{CLat}_{\sqcap}}(p,q)$, $\dot{F}\sqsubseteq[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}$ if and only if $\tau:\dot{F}\circ\mathbf{\Omega}\mathbin{\dot{\to}}\mathbf{\Pi}$ in the fibration $\mathrm{cod}_{\mathrm{Id}_{A},q}$2.2). ∎
@@
When the indexing set $A$ is one-point $1$, the codensity lifting $[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}$ is determined by the object $[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}\mathbf{\Omega}$.
@@
Assume $A=1$ in the setting of Definition 4.6. Let $(F,{\dot{F}}):p\to q$ be a fibered one-cell. If $R^{p,\mathbf{\Omega}}\circ L^{p,\mathbf{\Omega}}=\mathrm{id}$ and $\dot{F}$ preserves fibered meets, the following statements hold.
@@
The assumption $R^{p,\mathbf{\Omega}}\circ L^{p,\mathbf{\Omega}}=\mathrm{id}$ is mild enough for all examples of ${\mathbf{CLat}_{\sqcap}}$-fibrations in this paper. The first statement means $\dot{F}$ has the same property as $[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}$ written in Proposition 4.8, and it immediately yields the second statement. This proposition will be used to get a criterion for checking $[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}=\times\colon\mathbb{E}^{2}\to\mathbb{E}$ (Corollary 5.1).

### v2

@@ 3.2 Composing $F$-Coalgebras
Let $F\colon\mathbb{C}\to\mathbb{C}$ be an endofunctor over a category $\mathbb{C}$. We model an $N$-ary composition operation for $F$-coalgebras by a pair of
@@
    a functor $T\colon\mathbb{C}^{N}\to\mathbb{C}$ (called structure functor) describing how the operation transforms coalgebra carriers, and
@@
The second component induces the lifting ${T}_{\lambda}\colon(\mathbf{Coalg}(F))^{N}\to\mathbf{Coalg}(F)$ of $T$ along $(U_{F})^{N},U_{F}$ given by
${T}_{\lambda}\coloneq\lambda\circ T(c_{1},\cdots,c_{N}).$
@@
We introduce a binary one-step composition operation (×$(\times,\lambda)$ on ${\mathcal{P}}$-coalgebras. It takes the binary product of coalgebra carriers. The distributive law $\lambda^{{\mathcal{P}}}_{X,Y}\colon{\mathcal{P}}X\times{\mathcal{P}}Y\to{\mathcal{P}}(X\times Y)$ is given by $\lambda^{{\mathcal{P}}}_{X,Y}(A,B)\coloneq A\times B$.
@@
$\lambda^{\mathrm{da}}_{X,Y}((t_{1},\rho_{1}),(t_{2},\rho_{2}))\coloneq(t_{1}\land t_{2},~{}\lambda a~{}.~{}(\rho_{1}(a),\rho_{2}(a))).$
@@ 3.3 Composing Coalgebraic Bismulations
{mdframed}
Let $(F,\dot{F})\colon p\to p$ be a 1-cell in $\mathbf{CAT}^{\mathbf{CLat}_{\sqcap}}$. How can we lift ${T}_{\lambda}$ along $(\mathbf{Coalg}(p))^{N},\mathbf{Coalg}(p)$?
@@
Let $(F,{\dot{F}})\colon p\to p$ and $(T,{\dot{T}})\colon p^{N}\to p$ be 1-cells in $\mathbf{CAT}^{\mathbf{CLat}_{\sqcap}}$ and $(\lambda,\dot{\lambda})\colon(T,{\dot{T}})\circ(F^{N},{\dot{F}}^{N})\to(F,{\dot{F}})\circ(T,{\dot{T}})$ be a 2-cell. Then $(T_{\lambda},\dot{T}_{\dot{\lambda}})\colon(\mathbf{Coalg}(p))^{N}\to\mathbf{Coalg}(p)$ is a 1-cell in $\mathbf{CAT}^{\to}$.∎
A restricted version of this theorem is also available. Let $C$ be a subclass of objects in $\mathbb{E}$. If $\lambda$ is liftable on $C^{N}$, then we obtain a 1-cell $(T_{\lambda},\dot{T}_{\dot{\lambda}})\colon(\mathbf{Coalg}(p)|_{C})^{N}\to\mathbf{Coalg}(p)$; here $\mathbf{Coalg}(p)|_{C}$ is the restriction of $\mathbf{Coalg}(p)$ to the category of ${\dot{F}}$-coalgebras over objects in $C$. This restriction will be used after Proposition 7.3.
@@
In the setting of Theorem 3.9, for any $F$-coalgebra $c_{i}\colon X_{i}\to F(X_{i})$ ($i=1\cdots N$),
@@
{mdframed}
Let $(T,\lambda)$ be an $N$-ary one-step composition operation for $F$-coalgebras.
  1. 1.
    How do we non-trivially lift structure functors (such as $N$-ary product functors) in order to capture composition operators at the level of relations?
  2. 2.
    When is $\lambda$ liftable as a distributive law between codensity liftings of $F$ and $T$?
To answer to these questions, we employ Beohar et al.’s recent decomposition result of codensity liftings [3]. Their decomposition separates fibration-specific parts in the codensity lifting from a central part that does the actual lifting task. Upon their decomposition, in Section 5, we first address (1) by generalizing the codensity lifting to arbitrary functors. Then in Section 6, we address (2) by providing a sufficient condition to lift $\lambda$ to a distributive law between codensity liftings. The proof of the sufficiency takes advantage of the decomposition and a 2-categorical nature of the central part in the decomposition.
@@ 4 Generalizing the Codensity Lifting
We recall the multiple parameter codensity lifting in [29]. It lifts an endofunctor $F\colon\mathbb{B}\to\mathbb{B}$ along a ${\mathbf{CLat}_{\sqcap}}$-fibration $p\colon\mathbb{E}\to\mathbb{B}$ using three parameters $\Omega,\mathbf{\Omega},\tau$. We pack $(p,\Omega,\mathbf{\Omega})$ into the following data.
@@
Let $A$ be a discrete category. A ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values is a tuple $(p\colon\mathbb{E}\to\mathbb{B},\Omega\colon A\to\mathbb{B},\mathbf{\Omega}\colon A\to\mathbb{E})$ of functors such that $\Omega=p\circ\mathbf{\Omega}$ and $p$ is a ${\mathbf{CLat}_{\sqcap}}$-fibration. The notation for a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values is $(\Omega,\mathbf{\Omega})\colon Id_{A}\to(p\colon\mathbb{E}\to\mathbb{B})$.
Let $(\Omega,\mathbf{\Omega})\colon Id_{A}\to(p\colon\mathbb{E}\to\mathbb{B})$ be a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values. The codensity lifting of a functor $F\colon\mathbb{B}\to\mathbb{B}$ takes as argument a natural transformation $\tau\colon F\circ\Omega\to\Omega$. The components $\tau_{a}\colon F\circ\Omega(a)\to\Omega(a)$ are referred to in coalgebraic modal logic as modalities; $\tau$ is just an indexed collection of these, and below we will often refer to $\tau$ itself as a modality. The codensity lifting of $F\colon\mathbb{B}\to\mathbb{B}$ with $\tau$ along $p$ is given by
$[F,\tau]^{\mathbf{\Omega}}X\coloneq\bigsqcap_{a\in A,k\in\mathbb{E}(X,\mathbf{\Omega}(a))}(\tau_{a}\circ F(pk))^{*}(\mathbf{\Omega}(a)).$
It is an endofunctor over $\mathbb{E}$, and is a lifting of $F$ along $p$. When $A=1$, it is called the single-parameter codensity lifting. ${\dot{F}}$-bisimulations (resp. ${\dot{F}}$-bisimilarities) where ${\dot{F}}$ is a codensity lifting are called codensity bisimulations (resp. codensity bisimilarities).
@@
We write $2$ for the two-point set $\{\mathrm{false},\mathrm{true}\}$ and $\mathrm{Eq}_{2}$ for the object $(2,\{(x,x)~{}|~{}x\in 2\})$ in $\mathbf{EqRel}$. We identify it as a functor of type $1\to\mathbf{EqRel}$, and form a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $(2,\mathrm{Eq}_{2})\colon Id_{1}\to U_{\mathbf{EqRel}}$. The codensity lifting of ${\mathcal{P}}$ with the modality $\diamond\colon{\mathcal{P}}(2)\to 2$ given by $\diamond(U)=true\iff true\in U$ coincides with the lifting $\dot{\mathcal{P}}$ in Example 3.3 [23]. Therefore the standard bisimulations on ${\mathcal{P}}$-coalgebras are expressible as codensity bisimulations.
@@
We adopt the following ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values   $(\Omega,\mathbf{\Omega})\colon Id_{\Sigma\uplus\{\epsilon\}}\to U_{\mathbf{EqRel}}$ where $\Omega$ and $\mathbf{\Omega}$ are functors constantly returning $2$ and $\mathrm{Eq}_{2}$, respectively. We define the modality for ${F_{\mathrm{da}}}$ by
@@
Codensity bisimulation data consists of a ${\mathbf{CLat}_{\sqcap}}$-fibration with truth values $(\Omega,\mathbf{\Omega})\colon Id_{A}\to(p\colon\mathbb{E}\to\mathbb{B})$, a functor $F\colon\mathbb{B}\rightarrow\mathbb{B}$, and a natural transformation $\tau\colon F\circ\Omega\to\Omega$.
@@ 4.1 Decomposition of Codensity Lifting
Let $((\Omega,\mathbf{\Omega})\colon Id_{A}\to(p\colon\mathbb{E}\to\mathbb{B}),F,\tau)$ be codensity bisimulation data. Recently, Beohar et al. [3] introduced a decomposition of the codensity lifting as a “sandwich” of a monotone function within Galois connections. We adopt their decomposition as it is useful for analyzing the interaction between codensity liftings. In this paper, we describe their decomposition in fibered category theory. The decomposition is given as
@@
The description of the components of (2) is in order. First, we observe that the product $[A,U_{\mathbf{Pred}}^{op}]\colon[A,\mathbf{Pred}^{op}]\to[A,\mathbf{Set}^{op}]$ of $A$-fold copies of the opposite of the subobject fibration $U_{\mathbf{Pred}}\colon\mathbf{Pred}\to\mathbf{Set}$ is a ${\mathbf{CLat}_{\sqcap}}$-fibration. Now the adjunction $L^{p,\mathbf{\Omega}}\dashv R^{p,\mathbf{\Omega}}$ in (2) is given between $\mathbb{E}$ and the category obtained by the change-of-base of the fibration $[A,U_{\mathbf{Pred}}^{op}]$ along the functor $H(b)\coloneq(\mathbb{B}(b,\Omega(-)))^{op}$:
@@
Concretely, an object of $\mathrm{Sp}^{A}(\mathbb{B},\Omega)$ is a pair of an object $X\in\mathbb{B}$ and a mapping $S\colon A\to\mathbf{Set}$ such that for any $a\in A$, $S(a)\subseteq\mathbb{B}(X,\Omega(a))$. A morphism from $(X,S)$ to $(Y,T)$ is a morphism $f\colon X\to Y$ such that for any $a\in A$ and $k\in T(a)$, $k\circ f\in S(a)$. [^2] This category is similar to the one of topological spaces. Put $A=1$. The mapping $S$ specifies a subset of $\mathbb{B}(X,\Omega)$, which may be seen as a topology on $X$. An element in $S$ corresponds to an open subset, and the condition on morphism corresponds to the backword preservation of open subsets. For this reason we name this category Sp. The evident projection functor $r^{A}_{\mathbb{B},\Omega}\colon\mathrm{Sp}^{A}(\mathbb{B},\Omega)\to\mathbb{B}$ is a ${\mathbf{CLat}_{\sqcap}}$-fibration, because $s$ is so.
@@
$L^{p,\mathbf{\Omega}}(P)$ :-$\coloneq(p(P),\lambda a~{}.~{}p(\mathbb{E}(P,\mathbf{\Omega}(a)))),:$ $L^{p,\mathbf{\Omega}}(f)\coloneq f,$
@@
The middle part $\mathrm{Sp}^{A}(\mathbb{B},\Omega)$ in (2) is defined as a lifting of an endofunctor $F\colon\mathbb{B}\to\mathbb{B}$ along $r^{A}_{\mathbb{B},\Omega}$ using a modality $\tau\colon F\circ\Omega\to\Omega$. It is an extension of the predicate lifting $(P\colon X\to\Omega)\mapsto(\tau\circ F(P)\colon F(X)\to\Omega)$, which is commonly used in the theory of coalgebras, to sets of morphisms into $\Omega$.
$\mathrm{Sp}^{A}(F,\tau)(X,S)$ :-$\coloneq(FX,\lambda a~{}.~{}\{\tau_{a}\circ Fk~{}|~{}k\in S(a)\}):$ (3)
$\mathrm{Sp}^{A}(F,\tau)(f)$ :-$\coloneq Ff.:$ (4)
@@ 4.2 Generalizing Codensity Liftings
  • A 0-cell is a pair of a category $\mathbb{C}$ with a functor $\Omega\colon A\to\mathbb{C}$.
  • A 1-cell from $(\mathbb{C},\Omega)$ to $(\mathbb{D},\Pi)$ is a pair of a functor $F\colon\mathbb{C}\to\mathbb{D}$ and a natural transformation $\tau\colon F\circ\Omega\to\Pi$.
  • A 2-cell from $(F,\tau)\colon(\mathbb{C},\Omega)\to(\mathbb{D},\Pi)$ to $(G,\upsilon)$ is a natural transformation $\alpha\colon F\to G$ such that
Observe that an endo-1-cell on $(\mathbb{C},\Omega)$ bijectively corresponds to a pair of an endofunctor $F$ on $\mathbb{C}$ and a modality $\tau\colon F\circ\Omega\to\Omega$.
@@
Here, the functor $\mathrm{Sp}^{A}(F,\tau)$ is the extension of (3) to arbitrary 1-cell $(F,\tau)\colon(\mathbb{C},\Omega)\to(\mathbb{D},\Pi)$ by the same formulas (3),(4). Moreover, $\mathrm{Sp}^{A}(F,\tau)$ preserves Cartesian morphisms and all meets. ∎
@@
Let $(\Omega,\mathbf{\Omega})\colon Id_{A}\to(p\colon\mathbb{E}\to\mathbb{B})$ and $(\Pi,\mathbf{\Pi})\colon Id_{A}\to(q\colon\mathbb{F}\to\mathbb{C})$ be ${\mathbf{CLat}_{\sqcap}}$-fibrations with truth values. The codensity lifting of $F\colon\mathbb{B}\to\mathbb{C}$ along $p,q$ with a natural transformation $\tau\colon F\circ\Omega\to\Pi$, denoted by $[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}$, is defined by
$[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}$ :-$\coloneq R^{q,\mathbf{\Pi}}\circ\mathrm{Sp}^{A}(F,\tau)\circ L^{p,\mathbf{\Omega}}:$ (5)
@@
For any $(F,\dot{F})\in\mathbf{CAT}^{\mathbf{CLat}_{\sqcap}}(p,q)$, $\dot{F}\sqsubseteq[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}$ if and only if $\tau\colon\dot{F}\circ\mathbf{\Omega}\mathbin{\dot{\to}}\mathbf{\Pi}$ in the fibration $\mathrm{cod}_{\mathrm{Id}_{A},q}$2.2). ∎
@@
When the indexing set $A$ is one-point $1$, the codensity lifting $[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}$ is determined by the object $[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}\mathbf{\Omega}$. We can utilize it for general cases $(A\neq 1)$ because a codensity lifting is a meet $\sqcap_{a\in A}[F,\tau_{a}]^{\mathbf{\Omega}(a),\mathbf{\Pi}(a)}$ of $A$, and we have that $[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}$ is determined by objects $[F,\tau_{a}]^{\mathbf{\Omega}(a),\mathbf{\Pi}(a)}\mathbf{\Omega}(a)$.
@@
Assume $A=1$ in the setting of Definition 4.6. Let $(F,{\dot{F}})\colon p\to q$ be a fibered one-cell. If $R^{p,\mathbf{\Omega}}\circ L^{p,\mathbf{\Omega}}=\mathrm{id}$ and $\dot{F}$ preserves fibered meets, the following statements hold.
@@
The assumption $R^{p,\mathbf{\Omega}}\circ L^{p,\mathbf{\Omega}}=\mathrm{id}$ is mild enough for all examples of ${\mathbf{CLat}_{\sqcap}}$-fibrations in this paper. The first statement means $\dot{F}$ has the same property as $[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}$ written in Proposition 4.8, and it immediately yields the second statement. This proposition will be used to get a criterion for analysing $[F,\tau]^{\mathbf{\Omega},\mathbf{\Pi}}=\times\colon\mathbb{E}^{2}\to\mathbb{E}$ (Corollary 5.1).